40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Mar 4 10:54:12 2021, Paco, Anchal, Summary, LSC, POY11 measurement, tried to lock Green Yend laser 20210304_POY11_Spectrum_YARMLocked.pdf20210304_POY11_Spectrum_YARMLocked.tar.gz
    Reply  Thu Mar 4 11:59:25 2021, Paco, Anchal, Summary, LSC, Watchdog tripped, Optics damped back 
       Reply  Thu Mar 4 15:48:26 2021, Koji, Summary, PEM, Watchdog tripped, Optics damped back 
Message ID: 15863     Entry time: Thu Mar 4 15:48:26 2021     In reply to: 15862
Author: Koji 
Type: Summary 
Category: PEM 
Subject: Watchdog tripped, Optics damped back 

EQs seen on Summary pages
https://nodus.ligo.caltech.edu:30889/detcharsummary/day/20210304/pem/seismic_blrms/

ELOG V3.1.3-